OpenStudy (jv0408):

The table lists the values for two parameters, x and y, of an experiment. What is the approximate value of y for x = 4.0? x y 2.5 6.25 9.4 88.36 15.6 243.36 19.5 380.25 25.8 665.64

OpenStudy (zale101):

By looking at the plotted points, what is the rule in this function?

OpenStudy (jv0408):

im stuped on it give me a hint

OpenStudy (zale101):

What is the function/rule when you plug x=2.5 you will get 6.25?

OpenStudy (jv0408):

*3.5?

OpenStudy (zale101):

You mean 2.5 times 3.5 gives you 6.25?

OpenStudy (jv0408):

2*3=6

OpenStudy (jv0408):

5*5=25

OpenStudy (jv0408):

?

OpenStudy (zale101):

No, 2.5*3.5=8.75 not 6.25

OpenStudy (jv0408):

oh ok

OpenStudy (jv0408):

is it 2.5?

OpenStudy (zale101):

Yes.

OpenStudy (jv0408):

ok now what do I do?

OpenStudy (zale101):

if x=2.5, then x*x which means 2.5*2.5= 6.25. What is the rule? I already mentioned it.

OpenStudy (zale101):

Let us try 9.4. What rule when x=9.4 gives you 88.36?

OpenStudy (jv0408):

x=4.0

OpenStudy (zale101):

what is 9.4 * 4.0?

OpenStudy (jv0408):

no x=9.4

OpenStudy (zale101):

Yes.

OpenStudy (zale101):

What is the rule here? Can you write the equation for y now?

OpenStudy (jv0408):

y=x*x?